OpenStudy (anonymous):

How many solutions does the equation 3x − 7 = 4 + 6 + 4x have? Two None Infinitely many One

OpenStudy (anonymous):

do you know how to find x?

OpenStudy (anonymous):

Try simplifying the equation first: 3x -7 = 4x +10 Now try manipulate the constant like this: 3x = 4x +17 The rest is up to ya t o figure it out.

OpenStudy (anonymous):

what

OpenStudy (anonymous):

find x by moving all the terms that are the same to the other side. 3x - 7 = 4 + 6 + 4x So first add 7 to the right side of the equation side & subtract 4x to the left. what do you get?

OpenStudy (anonymous):

its 1 solution?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yep